# Tight Guarantees for Cut-Relative Survivable Network Design via a Decomposition Technique

### v2

@@ 6.1.1 Some remarks and discussion
The other direction is more tricky. Since $F\supseteq\bigcup_{i\in[k]}\delta_{G}(V_{i})$, we can focus on sets that are not a union of leaf-sets, and hence are not small-$(f,G)$-cuts. Recall that we say that two sets $A,B$ cross if $A-B,A\cap B$ and $B-A$ are all non-empty. Suppose $S\subseteq V$ crosses some $r$ leaf sets. Similar to the proof of Lemma 4.7, one potential approach would be to utilize the weak supermodularity of $\operatorname{def}_{f,F}$ to move to sets that cross fewer leaf sets, and thereby use induction on $r$ to establish that $\operatorname{def}_{f,F}(S)\leq 0$. There are some subtle issues that arise here that we want to highlight.
@@
The base cases, $r=0$ or $r=1$ are easy enough to argue, using the fact that $F(V_{i})$ is a feasible ${(\gamma_{i},G[{V_{i}}])}\text{-}\mathsf{NDP}$-solution for all $i\in[k]$. So suppose $r\geq 2$, A natural idea would be to pick some leaf set that $S$ crosses, say $V_{i}$, and apply weak supermodularity to $S$ and $V_{i}$, since we know that all four sets $S\cap V_{i},S\cup V_{i},S-V_{i},V_{i}-S$ cross fewer than $r$ leaf sets, and none of these sets is a union of leaf-sets and hence a small-$(f,G)$-cut. This would yield, for $\{X,Y\}=\{S\cap V_{i},S\cup V_{i}\}$ or $\{X,Y\}=\{S-V_{i},V_{i}-S\}$, that $\operatorname{def}_{f,F}(S)+\operatorname{def}_{f,F}(V_{i})\leq\operatorname{def}_{f,F}(X)+\operatorname{def}_{f,F}(Y)\leq 0$. However, since $V_{i}$ need not be a small-$(f,G)$-cut, this does not let us infer anything about $\operatorname{def}_{f,F}(S)$. To fix this, one could try to pick some small-$(f,G)$-cut $T\supseteq V_{i}$. But then we need not have that all the sets $S\cap T,S\cup T,S-T,T-S$ cross fewer than $r$ leaf sets. In particular, if $T$ contains all the same leaf-sets that $S$ crosses, then $S\cap T$ and $T-S$ would also cross all of these leaf sets.
@@
The key missing ingredient needed to make this approach work is the following leaf-set separation property: for any two leaf sets, there is some small-$(f,G)$-cut that separates them. This can be shown using insights from Lemma 6.1; we do not know of any other way of showing this. Given this property, one can complete the induction proof as follows. Let $V_{i}$ and $V_{j}$ be two leaf sets that $S$ crosses. Now let $T$ be a small-$(f,G)$-cut separating $V_{i}$ and $V_{j}$. Now, it is the case that all four sets $S\cap T,S\cup T,S-T,T-S$ cross fewer than $r$ leaf sets, so by weak-supermodularity and the induction hypothesis, we obtain that $\operatorname{def}_{f,F}(S)+\operatorname{def}_{f,F}(T)\leq 0$, which implies that $\operatorname{def}_{f,F}(S)\leq 0$, since $T$ is a small-$(f,G)$-cut.
To prove the leaf-set separation property, let $(f_{i},V_{i})$ and $(f_{j},V_{j})$ be two leaves of the canonical decomposition tree. Let $(h,S)$ be their least common ancestor, and let $(h_{1},A)$, $(h_{2},S-A)$ be the children of $(h,S)$. Clearly, $A$ separates $V_{i}$ and $V_{j}$. We have that $0<\operatorname{def}_{h,G[{S}]}(A)=\max\,\{\operatorname{def}_{h,G[{S}]}(T):T\subseteq S\}$. So for any $s\in A$, $t\in S-A$, by Lemma 6.1 (c), we may assume that $A=T\cap S$, where $\operatorname{def}_{f,G}(T)=\operatorname{def}_{h,G[{S}]}(A)$. It follows that $T$ is a small-$(f,G)$-cut that separates $V_{i}$ and $V_{j}$.
@@ 7 Hardness result
Consider the following path-relative $\mathsf{SNDP}$ instance. We add a source $s$ and sink $t$, and edges $sv,vt$ of cost 00, for all $v\in V$. Let $G^{\prime}=(V^{\prime},E^{\prime})$ denote this graph. The base-$\mathsf{SNDP}$ instance asks for $(k+n)$-edge connectivity between $s$ and $t$, where $n=|V|$. That is, the base cut-requirement function $f$ is given by $f(S^{\prime})=k+n$ if $S^{\prime}\subseteq V^{\prime}$ is an $s$-$t$ cut, and $f(S^{\prime})=0$ otherwise. We show that path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$ are actually equivalent in this case, and that they encode $k\text{-}\mathsf{ECSS}$ on the graph $G$.
@@
$\Rightarrow$ direction. Suppose that $H$ is feasible for path-relative $\mathsf{SNDP}$. If there is some edge $sv\notin H$, then consider the fault-set $F=\delta_{G^{\prime}}(s)-\{sv\}$. Then $H-F$ clearly has no $s$-$t$ path, since $\delta_{H-F}(s)=\emptyset$. But $G^{\prime}-F$ has an $s$-$t$ path $sv,vt$. So we must have $H\supseteq\delta_{G^{\prime}}(s)$. A similar argument shows that $H\supseteq\delta_{G^{\prime}}(t)$.
We next argue that $|\delta_{H}(S^{\prime})|\geq n+k$ for every $s$-$t$ cut $S^{\prime}$ for which $\emptyset\neq S:=S^{\prime}\cap V\subsetneq V$. This is equivalent to showing that $|\delta_{H(V)}(S)|\geq k$, since $|\delta_{H}(S^{\prime})\cap\delta_{G^{\prime}}(s)|+|\delta_{H}(S^{\prime})\cap\delta_{G^{\prime}}(t)|=n$. Assume that $s\in S^{\prime}$ without loss of generality. Suppose $|\delta_{H(V)}(S)|<k$. Since $|\delta_{G}(S)|\geq k$, there is some edge $uv\in\delta_{G}(S)-\delta_{H(V)}(S)$. Suppose $u\in S$. Consider the fault-set $F=\delta_{H(V)}(S)\cup\{sw:w\in V-S\}\cup\{wt:w\in S\}$. Then $|F|<k+n$ and $H-F$ has no $s$-$t$ path, since we have constructed $F$ to ensure that $\delta_{H-F}(S^{\prime})=\emptyset$. However, $G-F$ has an $s$-$t$ path: $su,uv,vt$. This contradicts that $H$ is feasible for path-relative $\mathsf{SNDP}$.
@@
Since the cost of the edges in $E^{\prime}-E$ is $0$, the above {cut, path}-relative $\mathsf{SNDP}$ instance is exactly the same as $k\text{-}\mathsf{ECSS}$ on the graph $G$.
@@ Appendix A Non-equivalence of path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$
We consider the same graph and base-SNDP problem as in Fig. 1. The graph is reproduced below for convenience, and recall that the base cut-requirement function models $s$-$t$ $2$-edge connectivity: so $f(S)=2$ if $S$ is an $s$-$t$ cut, and $f(S)=0$ otherwise.
@@
While cut-relative $\mathsf{SNDP}$ and path-relative $\mathsf{SNDP}$ are not in general equivalent, they are closely related. Dinitz et al. [9] showed (see Lemma A.1 in [9]) that one can give a cut-based formulation for path-relative $\mathsf{SNDP}$, where $H$ is feasible iff $|\delta_{H}(S)|\geq\min\bigl{\{}f(S),|\delta_{G}(S)|\}$ for every $S\subseteq V$ such that $G[{S}]$ and $G[{V^{\prime}-S}]$ are connected graphs, where $V^{\prime}\subseteq V$ is the vertex-set of the component containing $S$. That is, we require the cut-constraints of $\mathsf{CR}\text{-}\mathsf{SNDP}$ to hold for a suitable collection of node-sets (as opposed to all node-sets in $\mathsf{CR}\text{-}\mathsf{SNDP}$). This immediately implies that $\mathsf{CR}\text{-}\mathsf{SNDP}$ and path-relative $\mathsf{SNDP}$ are equivalent when the underlying graph is a (capacitated) complete graph. Also, as mentioned earlier, path-relative $k\text{-}\mathsf{ECSS}$ and cut-relative $k\text{-}\mathsf{ECSS}$ are equivalent.
@@
This was shown by [8]. We include a proof for completeness. We have already seen that a feasible solution to $\mathsf{CR}\text{-}\mathsf{SNDP}$ is also feasible for path-relative $\mathsf{SNDP}$. Now suppose that $H\subseteq E$ is a feasible solution to path-relative $k\text{-}\mathsf{ECSS}$. Suppose $\delta_{H}(S)<\min\bigl{\{}k,|\delta_{G}(S)|\bigr{\}}$ for some set $\emptyset\neq S\subsetneq V$. Consider $F=|\delta_{H}(S)|$, and let $e=uv\in\delta_{G-F}(S)$. Then, $|F|<k$ and $G-F$ has a $u$-$v$ path, but $H-F$ has no $u$-$v$ path, which contradicts that $H$ is feasible for path-relative $k\text{-}\mathsf{ECSS}$.
@@
where the second inequality is because $A$ is a maximum-deficiency set. If both $X$ and $Y$ are small-$(f,G)$-cuts, then we have $x_{e}=1$ for all $e\in\delta_{G}(X)\cup\delta_{G}(Y)\cup Z$, and hence $x_{e}=1$ for all $e\in\delta_{G}(T)$. Otherwise, $\operatorname{wdef}_{f,x}(T)+\operatorname{wdef}_{f,X}(A)\leq\operatorname{wdef}_{f,x}(X)+\operatorname{wdef}_{f,x}(Y)$, coupled with (9) shows that $\operatorname{wdef}_{f,x}(T)\leq\min\bigl{\{}\operatorname{wdef}_{f,x}(X),\operatorname{wdef}_{f,x}(Y)\bigr{\}}$, which is at most 00 since at least one of $X$, $Y$ is not a small-$(f,G)$-cut. This shows that $x$ is feasible to $(\text{{CRLP}}_{(f,G)})$. ∎

### v3

@@ 6.1.1 Further remarks and discussion
The other direction is more tricky. Any small-$(f,G)$-cut AA must be a union of leaf-sets, because if $A\cap V_{i},V_{i}-A$ are both non-empty for some leaf-set $V_{i}$, then Lemma 6.1 (c) implies that there is some small-$(f_{i},G[{V_{i}}])$-cut, which contradicts that $(f_{i},V_{i})$ is a leaf of the canonical decomposition tree. Since $F\supseteq\bigcup_{i\in[k]}\delta_{G}(V_{i})$, we can focus on sets that are not a union of leaf-sets, and hence are not small-$(f,G)$-cuts. Recall that we say that two sets $A,B$ cross if $A-B,A\cap B$ and BAB-A are all non-empty. Suppose $S\subseteq V$ crosses some $r$ leaf sets. Similar to the proof of Lemma 4.7, one potential approach would be to utilize the weak supermodularity of $\operatorname{def}_{f,F}$ to move to sets that cross fewer leaf sets, and thereby use induction on $r$ to establish that $\operatorname{def}_{f,F}(S)\leq 0$. There are some subtle issues that arise here.
@@
The base cases, $r=0$ or $r=1$ are easy enough to argue, using the fact that $F(V_{i})$ is a feasible ${(\gamma_{i},G[{V_{i}}])}\text{-}\mathsf{NDP}$-solution for all $i\in[k]$. So suppose $r\geq 2$, A natural idea would be to pick some leaf set that $S$ crosses, say $V_{i}$, and apply weak supermodularity to $S$ and $V_{i}$, since we know that all four sets $S\cap V_{i},S\cup V_{i},S-V_{i},V_{i}-S$ cross fewer than $r$ leaf sets, and none of these sets is a union of leaf-sets and hence a small-$(f,G)$-cut. This would yield, for $\{X,Y\}=\{S\cap V_{i},S\cup V_{i}\}$ or $\{X,Y\}=\{S-V_{i},V_{i}-S\}$, that $\operatorname{def}_{f,F}(S)+\operatorname{def}_{f,F}(V_{i})\leq\operatorname{def}_{f,F}(X)+\operatorname{def}_{f,F}(Y)\leq 0$. However, since $V_{i}$ need not be a small-$(f,G)$-cut, this does not let us infer anything about $\operatorname{def}_{f,F}(S)$. To fix this, we would like to pick some small-$(f,G)$-cut $T\supseteq V_{i}$, such that all the sets $S\cap T,S\cup T,S-T,T-S$ cross fewer than $r$ leaf sets (so that one can apply the induction hypothesis to these four sets). In particular, we need to avoid the situation where $T$ contains all the same leaf-sets that $S$ crosses, because then $S\cap T$ and TST-S would also cross all of these leaf sets.
@@
Fortunately, the proof of Lemma 6.6 shows that any two leaf-sets can be separated by some small-$(f,G)$-cut. (This utilizes insights from Lemma 6.1; we do not know of any other way of proving this.) This “leaf-set separation” property supplies the key missing ingredient needed to make the above induction proof work. Let $V_{i}$ and $V_{j}$ be two leaf sets that $S$ crosses. Now let $T$ be a small-$(f,G)$-cut separating $V_{i}$ and $V_{j}$. Now, all four sets $S\cap T,S\cup T,S-T,T-S$ cross fewer than $r$ leaf sets, so by weak-supermodularity and the induction hypothesis, we obtain that $\operatorname{def}_{f,F}(S)+\operatorname{def}_{f,F}(T)\leq 0$, which implies that $\operatorname{def}_{f,F}(S)\leq 0$, since $T$ is a small-$(f,G)$-cut.
Alternate decomposition.
The succinct description of the $\gamma_{i}$ functions in Lemma 6.8 suggests another decomposition into ${f}\text{-}\mathsf{NDP}$ instances, where the cut-requirement functions are now defined directly on the components of GZG-Z in a manner similar to the $\gamma_{i}$ functions. Recall that $C_{1},\ldots,C_{p}$ are the components of GZG-Z. To elaborate, for $j\in[p]$, define $\kappa_{j}:2^{C_{j}}\mapsto\mathbb{Z}$ as follows: $\kappa_{j}(\emptyset)=\kappa_{j}(C_{j}):=0$, and
$\kappa_{j}(S):=\max\,\Bigl{\{}f(S\cup X)-|\delta_{Z}(S\cup X)|:X=\bigcup_{j^{\prime}\in I}C_{j^{\prime}},\ I\subseteq[p]-\{j\}\Bigr{\}}\quad\forall\emptyset\neq S\subsetneq C_{j}.$
We can show that the $(\kappa_{1},C_{1}),\ldots,(\kappa_{p},C_{p})$ tuples satisfy the properties stated in Theorem 4.1. Consider a component $C_{j}$ and leaf-set $V_{i}$ containing $C_{j}$. Let $\mathsf{EQ}_{i}\subseteq[p]$ be such that $V_{i}=\bigcup_{j^{\prime}\in\mathsf{EQ}_{i}}C_{j^{\prime}}$. To gain intuition, consider any $S\subseteq C_{j}$ and consider the contraints imposed by $\gamma_{i}$ on $\delta_{G[{V_{i}}]}(S)$. Observe that $\delta_{G[{V_{i}}]}(C_{j^{\prime}})=\emptyset$ for any $j^{\prime}\in\mathsf{EQ}_{i}$, since $\delta_{G}(C_{j^{\prime}})\subseteq Z$ and there are no edges of $Z$ internal to $V_{i}$. So for any $Y=\bigcup_{j^{\prime}\in J}C_{j^{\prime}}$, where $J\subseteq\mathsf{EQ}_{i}-\{j\}$, we can infer that $\gamma_{i}$ imposes that we pick $\gamma_{i}(S\cup Y)$ edges from $\delta_{G[{C_{j}}]}(S)$. We argue that $\kappa_{j}$ is simply the function that aggregates these constraints. More precisely, we show that
$\kappa_{j}(S)=\max\,\Bigl{\{}\gamma_{i}(S\cup Y):Y=\bigcup_{j^{\prime}\in J}C_{j^{\prime}},\ J\subseteq\mathsf{EQ}_{i}-\{j\}\Bigr{\}}\quad\forall\emptyset\neq S\subsetneq C_{j}.$ (9)
This also implies that the $(\kappa_{j},C_{j})$ tuples satisfy the properties in Theorem 4.1.
As in the proof of Lemma 6.8, it is easy to see that for any $S\subseteq C_{j}$, we have $\kappa_{j}(S)$ is at least the RHS of (9). Consider any $Y=\bigcup_{j^{\prime}\in J}C_{j^{\prime}}$, $J\subseteq\mathsf{EQ}_{i}-\{j\}$. (Recall that $\mathcal{C}_{i}=\bigl{\{}\bigcup_{j\in I}V_{j}:I\subseteq[k]-\{i\}\bigr{\}}$, where $V_{1},\ldots,V_{k}$ are the leaf sets of the canonical decomposition.) If $\gamma_{i}(S\cup Y)=f(S\cup Y\cup X)-|\delta_{Z}(S\cup Y\cup X)|$, where $X\in\mathcal{C}_{i}$, then we can see that $Y\cup X$ is of the form $\bigcup_{j^{\prime}\in I}C_{j^{\prime}}$ for some $I\subseteq[p]-\{j\}$, and so $\kappa_{j}(S)\geq f(S\cup Y\cup X)-|\delta_{Z}(S\cup Y\cup X)|=\gamma_{i}(S\cup Y)$. To show that $\kappa_{j}(S)$ is at most the RHS of (9), suppose $\kappa_{j}(S)=f(S\cup X)-|\delta_{Z}(S\cup X)|$, where $X=\bigcup_{j^{\prime}\in I}C_{j^{\prime}}$, $I\subseteq[p]-\{j\}$. Let $T=S\cup X$. So $\operatorname{def}_{\kappa_{j},G[{V_{i}}]}(S)=\operatorname{def}_{f,G}(T)$. As in the proof of Lemma 6.8, by Lemma 6.1 (c), we have $\operatorname{def}_{f,G}(T)\leq\operatorname{def}_{f_{i},G[{V_{i}}]}(T\cap V_{i})$. But $T\cap V_{i}$ is of the form $S\cup Y$, where $Y=\bigcup_{j^{\prime}\in J}C_{j^{\prime}}$, $J\subseteq\mathsf{EQ}_{i}-\{j\}$, so we obtain that $\operatorname{def}_{\kappa_{j},G[{V_{i}}]}(S)\leq\operatorname{def}_{f_{i},G[{V_{i}}]}(S\cup Y)$, which translates to $\kappa_{j}(S)\leq f_{i}(S\cup Y)=\gamma_{i}(S\cup Y)$.
@@ 7 Hardness result
Consider the following path-relative $\mathsf{SNDP}$ instance. We add a source $s$ and sink $t$, and edges $sv,vt$ of cost 0, for all $v\in V$. Let $G^{\prime}=(V^{\prime},E^{\prime})$ denote this graph. The base-$\mathsf{SNDP}$ instance asks for $(k+n)$-edge connectivity between $s$ and $t$, where $n=|V|$. That is, the base cut-requirement function $f$ is given by $f(S^{\prime})=k+n$ if $S^{\prime}\subseteq V^{\prime}$ is an $s$-$t$ cut, and $f(S^{\prime})=0$ otherwise. We show that path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$ are actually equivalent in this case, and that they encode $k\text{-}\mathsf{ECSS}$ on the graph $G$.
@@
[…]
We next argue that $|\delta_{H}(S^{\prime})|\geq n+k$ for every $s$-$t$ cut $S^{\prime}$ for which $\emptyset\neq S:=S^{\prime}\cap V\subsetneq V$. This is equivalent to showing that $|\delta_{H(V)}(S)|\geq k$, since $|\delta_{H}(S^{\prime})\cap\delta_{G^{\prime}}(s)|+|\delta_{H}(S^{\prime})\cap\delta_{G^{\prime}}(t)|=n$. Assume that $s\in S^{\prime}$ without loss of generality. Suppose $|\delta_{H(V)}(S)|<k$. Since $|\delta_{G}(S)|\geq k$, there is some edge $uv\in\delta_{G}(S)-\delta_{H(V)}(S)$. Suppose $u\in S$. Consider the fault-set $F=\delta_{H(V)}(S)\cup\{sw:w\in V-S\}\cup\{wt:w\in S\}$. Then $|F|<k+n$ and HFH-F has no $s$-$t$ path, since we have constructed $F$ to ensure that $\delta_{H-F}(S^{\prime})=\emptyset$. However, GFG-F has an $s$-$t$ path: $su,uv,vt$. This contradicts that $H$ is feasible for path-relative $\mathsf{SNDP}$.
@@
[…]
@@ Appendix A Non-equivalence of path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$
We consider the same graph and base-SNDP problem as in Fig. 1. The graph is reproduced below for convenience, and recall that the base cut-requirement function models $s$-$t$ 22-edge connectivity: so $f(S)=2$ if $S$ is an $s$-$t$ cut, and $f(S)=0$ otherwise.
@@
While cut-relative $\mathsf{SNDP}$ and path-relative $\mathsf{SNDP}$ are not in general equivalent, they are closely related. Dinitz et al. [9] showed (see Lemma A.1 in [9]) that one can give a cut-based formulation for path-relative $\mathsf{SNDP}$, where $H$ is feasible iff $|\delta_{H}(S)|\geq\min\bigl{\{}f(S),|\delta_{G}(S)|\}$ for every $S\subseteq V$ such that G[S]G[{S}] and $G[{V^{\prime}-S}]$ are connected graphs, where $V^{\prime}\subseteq V$ is the vertex-set of the component containing $S$. That is, we require the cut-constraints of $\mathsf{CR}\text{-}\mathsf{SNDP}$ to hold for a suitable collection of node-sets (as opposed to all node-sets in $\mathsf{CR}\text{-}\mathsf{SNDP}$). This immediately implies that $\mathsf{CR}\text{-}\mathsf{SNDP}$ and path-relative $\mathsf{SNDP}$ are equivalent when the underlying graph is a (capacitated) complete graph. Also, as mentioned earlier, path-relative $k\text{-}\mathsf{ECSS}$ and cut-relative $k\text{-}\mathsf{ECSS}$ are equivalent.
@@
This was shown by [8]. We include a proof for completeness. We have already seen that a feasible solution to $\mathsf{CR}\text{-}\mathsf{SNDP}$ is also feasible for path-relative $\mathsf{SNDP}$. Now suppose that $H\subseteq E$ is a feasible solution to path-relative $k\text{-}\mathsf{ECSS}$. Suppose $\delta_{H}(S)<\min\bigl{\{}k,|\delta_{G}(S)|\bigr{\}}$ for some set $\emptyset\neq S\subsetneq V$. Consider $F=|\delta_{H}(S)|$, and let $e=uv\in\delta_{G-F}(S)$. Then, $|F|<k$ and GFG-F has a $u$-$v$ path, but HFH-F has no $u$-$v$ path, which contradicts that $H$ is feasible for path-relative $k\text{-}\mathsf{ECSS}$.
@@
where the second inequality is because $A$ is a maximum-deficiency set. If both $X$ and $Y$ are small-$(f,G)$-cuts, then we have $x_{e}=1$ for all $e\in\delta_{G}(X)\cup\delta_{G}(Y)\cup Z$, and hence $x_{e}=1$ for all $e\in\delta_{G}(T)$. Otherwise, $\operatorname{wdef}_{f,x}(T)+\operatorname{wdef}_{f,X}(A)\leq\operatorname{wdef}_{f,x}(X)+\operatorname{wdef}_{f,x}(Y)$, coupled with (10) shows that $\operatorname{wdef}_{f,x}(T)\leq\min\bigl{\{}\operatorname{wdef}_{f,x}(X),\operatorname{wdef}_{f,x}(Y)\bigr{\}}$, which is at most 0 since at least one of $X$, $Y$ is not a small-$(f,G)$-cut. This shows that $x$ is feasible to $(\text{{CRLP}}_{(f,G)})$. ∎